Tuesday, 25 August

Tuesday, 25 August

Plenary Lectures

9:00 – 10:00  Auditorium Maximum H 105

George C. Papanicolaou, Stanford University, Stanford, USA Mathematical Problems in Geophysical Wave Propagation (I: p.17)

Chair: John Ball, Oxford University, Oxford, UK

10:15 – 11:15  Auditorium Maximum H 105

Tetsuji Miwa, RIMS, Kyoto University, Kyoto, Japan Algebraic Analysis of Solvable Lattice Models (I: p.16)

Chair: Yuri Manin, Max-Planck-Institut für Mathematik, Bonn, Germany

11:45 – 12:45  Auditorium Maximum H 105

Gilles Pisier, Université Paris 6, Paris, France and Texas A&M University, College Station, USA Operator Spaces and Similarity Problems (I: p.18)

Chair: Joram Lindenstrauss, Hebrew University, Jerusalem, Israel

Invited Lectures

\underline{14:00 – 14:45}

03 Loïc Merel, Université Paris 7 - Denis Diderot, Paris, France Determination of rational points using Dirichlet series (I: p.52) Chair: Wilhelm Zink

MA 005

08 Nicole Tomczak-Jaegermann, University of Alberta, Edmonton, Canada From finite- to infinite-dimensional phenomena in Geometric Functional Analysis, on local and asymptotic levels (I: p.108) Chair: W. Timothy Gowers

MA 001

12 Alain-Sol Sznitman, ETH Zürich, Zürich, Switzerland Brownian motion and random obstacles (I: p.149) Chair: Terry J. Lyons

H 1028

15 Percy Alec Deift, Courant Institute, New York University, New York, USA Uniform asymptotics for orthogonal polynomials (I: p.179) Chair: Tom H. Koornwinder

H 1012

16 Andreas W.M. Dress, Universität Bielefeld, Bielefeld, Germany The Tree of Life and Other Affine Buildings (I: p.186) Chair: Hanfried Lenz

H 2032

18 Mariolina Bartolini Bussi, Università di Modena, Modena, Italy Drawing instruments: the theory of a practice from history to didactics (I: p.197) Chair: Klaus-D. Graf

H 104

\underline{15:00 – 15:45}

!!!

Lecture by a Fields Medalist

— to be announced — Chair: N.N.

MA 001

03 Shinichi Mochizuki, RIMS, Kyoto University, Kyoto, Japan The Intrinsic Hodge Theory of p-adic Hyperbolic Curves (I: p.53) Chair: Helmut Koch

MA 005

12 Peter Gavin Hall, Australian National University, Canberra, Australia Applications of Intentionally Biased Bootstrap Methods (I: p.144) Chair: Olaf Bunke H 1028 15 Jan-Olov Strömberg, Royal Institute of Technology, Stockholm, Sweden Apriori estimates for approximation by wavelets in higher dimensions (I: p.180) Chair: Peter Maaß

H 1012

16 Bonnie Berger, Massachusetts Institute of Technology, Cambridge, USA Sequence-based Methods for Protein Folding (I: p.185) Chair: Michiel Hazewinkel

H 2032

18 James Stigler, University of California, Los Angeles, USA Classroom Mathematics Instruction in Three Cultures: An Introduction to the TIMSS Video Study (I: p.200) Chair: Thomas Jahnke H 104 \underline{16:15 – 17:00}

01 Matthew Foreman, University of California, Irvine, USA Generic Large Cardinals: New Axioms for Mathematics? (I: p.30) Chair: Peter G. Koepke

H 1012

02 Aner Shalev, The Hebrew University, Jerusalem, Israel Simple groups, permutation groups, and probability (I: p.45) Chair: Bernhard H. Neumann

MA 005

05 Boris Dubrovin, International School for Advanced Studies, Trieste, Italy and Steklov Math. Institute, Moscow, Russia Geometry and analytic theory of Frobenius manifolds (I: p.69) Chair: Alexander I. Bobenko

MA 001

11 Eugene Bogomolny, Université Paris Sud, Orsay, France Spectral Statistics (I: p.129) Chair: Pierre Collet

H 1028

13 Neil J.A. Sloane, AT&T Research Labs, Florham Park, USA Recent developments on the sphere-packing problem (I: p.161) Chair: Martin Aigner

H 2032

\underline{16:15 – 18:00}

18 George Eyre Andrews, Pennsylvania State University, University Park, USA D.J. Lewis (moderator), NSF, Arlington, USA David Alexander Smith, Duke University, Durham, USA Collegiate Mathematics Instruction for the 21st Century — A Panel Discussion (I: p.200) Chair:

H 104

\underline{17:15 – 18:00}

01 Ludomir Newelski, University of Wrocław, Wrocław and Polish Academy of Sciences, Poland Meager forking and m-independence (I: p.33) Chair: Alex James Wilkie

H 1012

02 Jeremy C. Rickard, University of Bristol, Bristol, England The abelian defect group conjecture (I: p.45) Chair: Dieter Happel

MA 005

11 Horng-Tzer Yau, Courant Institute, New York University, New York, USA Scaling Limit of Particle Systems, Incompressible Navier-Stokes Equation, and Boltzmann Equation (I: p.138) Chair: Ruedi Seiler

H 1028

13 Joseph Adolphe François Thas, University of Gent, Gent, Belgium Finite Geometries, Varieties and Codes (I: p.161) Chair: Theo Grundhöfer

H 2032

16 Rolf Rannacher, Universität Heidelberg, Heidelberg, Germany Mathematics in Numerical Flow Simulation (I: p.192) Chair: Rolf Jeltsch MA 001

Short Communications and Poster Sessions

Short Communications: Tuesday, 25 August, 14:00–14:45

14:00–14:15  14:15–14:30  14:30–14:45

Sect.: 01

Set Theory and Recursion Theory, Chair: Robert Fittler

MA 042 Philip D. Welch The Length of Infinite Time Turing Machine Computations (p.11)

Martin Zeman

◇ at Mahlo Cardinals (p.11)

Marat-M. Arslanov

The enumeration degree structures (p.5)

Sect.: 02

Rings and Modules II, Chair: Christine Bessenrodt

MA 311 Sergey P. Strunkov On p-blocks of minimal defect in a finite group and Gelfand pairs (p.29)

Johannes H. Meyer

Forcing Linearity Numbers (p.23)

Hossein Zand

Duality in radicals of associative rings (p.32)

Sect.: 05

Affine Differential Geometry, Chair: Salvador D.R. Gigena

MA 041 Kurt Leichtweiß On the infinitesimal affine rigidity of ellipsoids in the n-space (p.74)

Christine Scharlach

Affine Hyperspheres (p.79)

Salvador D.R. Gigena

On Affine Hypersurfaces with Constant Sectional Curvature (p.70)

Sect.: 05

General Geometric Structures on Manifolds, Chair: Lutz Weimann

MA 043 Sergey V. Zuev Analytic continuations of SU(2n)- invariant Ricci-flat Kähler metric Claudio Gorodski Contact Sub-Riemannian Symmetric Spaces (p.70)

Artur E. Fischer

The Adjoint Formulation and Symplectic Reduction (p.68)

Sect.: 06

2-Manifolds and Related Questions, Chair: Carl-Friedrich Bödigheimer

H 112 José-Luis Estévez A presentation for the mapping class group of a non-rientable surface from Hatcher and Thurston’s (p.87)

Dmitri V. Millionshikov

Krichever-Novikov algebras (p.93)

Andreas Zastrow

Planar Sets are Aspherical (p.100)

Sect.: 09

Large-Time Behavior, Chair: Martin Brokate

H 111 Ivan Kiguradze On explosive kneser solutions of nonlinear ordinary differential equations (p.179)

Matthias Wolfrum

A geometrical description for the attractors of semilinear parabolic PDE (p.194)

Martin Brokate

On a rate independent variational inequality (p.173)

Sect.: 09

Topological Dynamics, Chair: Fritz Boese

H 110 Sergey Bezuglyi Full groups of minimal homeomorphisms of Cantor sets (p.171)

Fritz Boese

Unilluminable Rooms (p.171)

Sect.: 11

q-Deformation, Quantization, Chair: Markus J. Pflaum

H 107 Hitoshi Konno Elliptic quantum groups (p.237)

Andreas Ruffing

On Schrödinger-Hermite Operators in Lattice Quantum Mechanics (p.243)

Markus J. Pflaum

Remarks on the quantization of stratified symplectic manifolds (p.242)

Sect.: 11

Mathematical Methods in Physics I, Chair: Mikhail Brodsky

H 1029 Boris V. Loginov Bifurcation theory methods in the problem about capillary-gravity waves in spatial layer of floating fluid (p.239)

Nina F. Golovanova

Eiconal expansion solution for T-matrix equation and the unitarity condition

(p.234)

Anton A. Golovanov

Solution of the T-matrixs equation by the eikonal expansion method (p.234)

Sect.: 13

Subgraphs, Chair: Béla Bollobás

H 106 Oliver M Riordan Spanning subgraphs of random graphs

(p.284)

Diane M. Spresser

Graphs with degree sets of maximal cardinality II (p.285)

Andrew G. Thomason

Hereditary hypergraph properties (p.286)

Sect.: 17

Distributed Parameter Systems II, Chair: Mikhail Zelikin

H 2033 Andrzej Just A Non-linear Hyperbolic Control Problem with Dirichlet Boundary Condition — Convergence of Galerkin Approximation

(p.343)

Dan Tiba

Recent Progress in the Analysis and Optimization of Plates (p.352)

Mikhail Zelikin

Riccati-type PDE and Multidimensional Calculus of Variations (p.356)

Sect.: 19

19th Century, Chair: Herbert A. David

H 2013 Mircea Radu Justus Grassmann’s Account of Synthesis in Mathematics: between Schelling and Kant (p.368)

Heiner Lichtenberg

The decoding of the Callippian Cycle in the Gregorian Calendar for the period between 1900 and 2099 (p.367)

Herbert A. David

Early sample measures of variability (p.366)

Short Communications: Tuesday, 25 August, 15:00–15:45

15:00–15:15  15:15–15:30  15:30–15:45

Sect.: 01

Model Theory and Recursion Theory, Chair: Andreas Baudisch

MA 042 Mikhail Peretyat’kin Expressive possibilities of finitely axiomatizable theories (p.9)

Niandong Shi

Universal graph problem and model theory (p.10)

Helmut Schwichtenberg

Ramification, Modality, and Linearity in Higher Type Recursion (p.10)

Sect.: 02

Miscellaneous, Chair: Corinna Wiedorn

MA 311 Martin Goldstern Interpolation of monotone functions in lattices (p.18)

Ming-Chang Kang

Rationality of Some Projective Linear Actions (p.19)

Vyacheslav A. Artamonov

Quantum polynomial algebra is a multiparameter deformation of an algebra of polynomials, i.e. it is the algebra of regular functions on a quantum affine space

(p.12)

Sect.: 05

Geodesics and Distance, Chair: Jin-ichi Itoh

MA 041 Minoru M.T. Tanaka Sard Theorem for the distance function

(p.82)

Lilia Del Riego

Exponential map for general sprays and connections (p.66)

Jin-ichi Itoh

The Lipschitz continuity of the distance function to the cut locus (p.72)

Sect.: 05

Lorentz Manifolds, Chair: Franki Dillen

MA 043 Andrzej Krolak Cauchy Horizon Endpoints and Differentiability (p.74)

Angel L. Garrido Bullon

Foliations of Lorentz manifolds (p.66)

Franki Dillen

Ruled Weingarten surfaces in Minkowski space (p.66)

Sect.: 06

Algebraic Topology III, Chair: Leonard Mdzinarishvili

H 112 Ivan Ivansic Borsuk’s Index and Pointed Movability for Projective Movable Continua (p.89)

Sibe Mardesic

Strong shape for topological spaces (p.92)

Mihkail M. Postnikov

Totality versus naturality (p.95)

Sect.: 08

Miscellaneous, Chair: Eberhard Kirchberg

MA 309 Eberhard Kirchberg Classification of non-simple nuclear C*- algebras Volodymyr K. Maslyuchenko Separately continuous mappings: Van Vleck spaces and inverse problems

Sect.: 09

Holomorphic and Symbolic Dynamics, Chair: Mitsuhiro Shishikura

H 111 Dierk Schleicher On iterated exponential maps (p.190)

George Osipenko

Applied Symbolic Dynamics (p.185)

Mitsuhiro Shishikura

Rigidity of complex quadratic polynomials and the universal Teichmüller space

(p.191)

Sect.: 11

Integrable Systems II, Chair: Karl-Georg Schlesinger

H 107 Vitaly Tarasov q-Hypergeometric solutions of the quantized Knizhnik-Zamolodchikov equation

(p.245)

Vladimir Dragovic

Integrable generalizations of classical mechanical systems (p.232)

Karl-Georg Schlesinger

Quantum geometry by non-abelian cohomology (p.244)

Sect.: 11

Mathematical Methods in Physics II, Chair: Boris V. Loginov

H 1029 Mikhail Brodsky A Physicist’s Definition of the Schwarz Function (p.229)

Joyati Debnath

Associated Transforms Through a Mathematical Software System (p.231)

Andrei Khrennikov

p-adic realization of negative probabilities and Einstein-Podolsky-Rosen paradox

(p.236)

Sect.: 13

Finite Groups and Cayley Graphs, Chair: Andreas W.M. Dress

H 106 Reinaldo Giudici Cayley’s Graphs of order 12 (p.276)

Cai-Heng Li

On Isomorphisms of Finite Cayley Graphs (p.280)

Mikhail H. Klin

On transitive permutation groups without semi-regular subgroups (p.279)

Sect.: 17

Stochastic Programming I, Chair: Werner Römisch

H 2033 Vladimir Norkin Stochastic optimization of discontinuous functions (p.346)

Nicole Groewe-Kuska

Confidence intervals for optimal values of simple recourse models (p.341)

Rüdiger Schultz

Mixed-Integer Linear Programming under Uncertainty (p.350)

Sect.: 19

Data Bases, Chair: Patrick D.F. Ion

H 2013 June E. Barrow-Green Mathematics in Britain 1860-1940: a database for the World Wide Web (p.365)

Harry B. Coonce

The Genealogy Project for Mathematicians (p.365)

Patrick D.F. Ion

The symbolism of mathematics: from pasigraphy in 1897 to MathML in 1998

(p.367)

Short Communications: Tuesday, 25 August, 16:15–17:00

16:15–16:30  16:30–16:45  16:45–17:00

Sect.: 03

Arithmetic Functions, Chair: Yasutoshi Nomura

MA 311 Philippe Ph. Elbaz-Vincent On Poly(ana)logs (p.35)

Arne Winterhof

On Incomplete Character Sums (p.47)

Yasutoshi Nomura

Arithmetic functions associated with arithmetic-geometric mean quotients

(p.44)

Sect.: 04

Surfaces/Higher-Dimensional Varieties II, Chair: Vasile P. Brinzǎnescu

MA 042 Alberto Conte On the unirationality of quartic hypersurfaces (p.52)

Fabrizio M.E. Catanese

Bidouble covers and surface geography (p.51)

Remo Gattazzo

Almost factorial quartic surfaces in ℙ_{ℂ}^{3} with a tacnodal point (p.52)

Sect.: 07

Various Topics I, Chair: Michael Forger

H 112 Lizhen Ji The trace class conjecture in the Selberg trace formula (p.106)

Alexander Baranov

Finitary simple Lie algebras (p.101)

Michael Forger

Symmetry and Symmetry Breaking: An Algebraic Approach to the Genetic Code

(p.104)

Sect.: 08

Hilbert Spaces, Wavelets, and Fractals, Chair: Gerald S. Goodman

H 110 Kofi V.S. Assiamoua A product of sesquilinear mappings on Hilbert spaces (p.114)

Raquel-Garcia Catalan

From MRA to Tight Frames (p.120)

Gerald S. Goodman

Dilation equations and Hutchinson fractals (p.126)

Sect.: 08

Real and Complex Approximation, Chair: Mirela Kohr

MA 041 Nikolai Ya. Radyno Some elements of a General Spectral Theory of Banach modules and their applications (p.150)

Oxana V. Motornaya

On the Approximation of the Functions by Algebraic Polynomials (p.142)

Andrey Olegovich Matveev

Error estimates for scattered data interpolation

Sect.: 08

Spectral Analysis and Operator Algebras, Chair: Sa Ge Lee

MA 309 Antoine-A.P. Bredimas Schlicht functions and exotic sprectral analysis: Applications to some conjectures and problems in analysis (p.118)

Valentyn Ja. Golodets

Dynamical entropy and quantum Ksystems (p.126)

Sa Ge Lee

The generalized normal state space and unital normal completely positive map

(p.139)

Sect.: 08

Matrix and Operator Functions, Chair: George N. Khimshiashvili

MA 043 Marcin J. Zygmunt Matrix Moment Problem (p.167)

Frank-Olme Speck

Relations between convolution type operators on intervals and on the half-line

(p.157)

George N. Khimshiashvili

Homotopy classes of elliptic boundary problems (p.134)

Sect.: 10

Elliptic Equations V (Fourier and Harmonic Analysis), Chair: Pawel Strzelecki

H 111 Alip Mohammed Riemann problem related to the Shilov boundary (p.214)

Dian K. Palagachev

Oblique derivative problem for uniformly elliptic operators with VMO coefficients and applications (p.216)

Paweł Strzelecki

Regularity for subelliptic p-harmonic maps and the ghost of Hardy spaces (p.222)

Sect.: 12

Brownian Motion I, Chair: Jean-Pierre Imhof

H 107 Frank B. Knight On the moments of the area functional of Brownian bridge (p.256)

Andrei Nikolaevich Borodin

Distributions of functionals of a particular non-Markov process (p.249)

Peter Mörters

Pathwise Kallianpur Robbins laws for planar Brownian motion (p.262)

Sect.: 12

Time Series Analysis, Chair: Alexander Frolovich Taraskin

H 1029 Mirta Bens̆ić Asymptotic properties of the least square estimators in regression models with singular errors (p.248)

Uwe Hassler

When Do Polynomial Regressions of Nonstationary Fractionally Integrated Processes Make Sense? (p.252)

Pavel S. Knopov

Asymptotic Properties of Non-Linear Estimators for Some Non-Stationary Regression Models (p.256)

Sect.: 15

Partial Differential Equations VI, Chair: Ivanka P. Nikolova

H 2033 Ivanka P. Nikolova Modeling of Fluid Flows with Strong Non-Linearity (p.307)

Christian von Törne

Method of Transport — A Multidimensional Scheme for Nonlinear Hyperbolic Systems (p.317)

Nicolai von Schroeders

Die numerische Lösung der schwachen Formulierung des Stokes-Systems im Außenraumgebiet (p.317)

Sect.: 16

Applications in Mechanics, Chair: Vadim D. Potapov

H 106 Mikhail Dmitrievich Kovalev Stability of hinged schemes and frameworks (p.326)

Xing Li

Doubly Periodic Punctured Plane Elasticity Problem with Relative Displacement

(p.327)

Vadim D. Potapov

Stability of systems under stochastic parametric excitation (p.331)

Sect.: 17

Game Theory, Chair: Hans Josef Pesch

H 2013 Arkadii A. Chikrii

— to be announced — (p.338)

Greta Tsolakovna Chikrii

Linear Differential Games with Variable Delay of Information (p.337)

Victor V. Zakharov

Subcore and reduced cooperative games (p.355)

Short Communications: Tuesday, 25 August, 17:15–18:00

17:15–17:30  17:30–17:45  17:45–18:00

Sect.: 03

Analytic Number Theory II, Chair: Jan-Christoph Puchta

MA 311 Julio Alcantara-Bode Reorderings of some divergent series and applications (p.34)

Michel L. Lapidus

Complex Dimensions of Fractal Strings and Oscillatory Phenomena in Fractal Geometry and Arithmetic (p.40)

Jan-Christoph Puchta

Estimating Dirichlet L-serieses close to the line 1 + it (p.45)

Sect.: 04

Surfaces/Higher-Dimensional Varieties III, Chair: Fabrizio M.E. Catanese

MA 042 Marina R. Marchisio Some new examples of smooth unirational quartic threefolds (p.56)

Uwe Nagel

Characterization of projective subschemes by locally free resolutions (p.57)

Tomasz Szemberg

Higher order embeddings and Seshadri constants of abelian varieties (p.61)

Sect.: 07

Various Topics II, Chair: Alexander Baranov

H 112 Thomas Bouetou About Bol algebras (p.102)

Lyudmyla L. Barannyk

On an algorithm for classifying subalgebras of the Poincaré algebra AP (2, n)

Sect.: 08

Analytic Functions II, Chair: Mohamed Abdel-Megied

H 110 Gueorgui S. Litvinchuk On solvability theory of Hilbert boundary value problem with a Carleman shift

(p.139)

Mirela Kohr

A direct boundary integral formulation for a Stokes flow problem (p.135)

Rauno Aulaskari

Q_{p}-spaces on Riemann surfaces (p.115)

Sect.: 08

Nonlinear Operators, Chair: Richard M. Dudley

MA 041 Radoje Šćepanović On Solvability of Some Nonlinear Equations in Hilbert Spaces (p.154)

Vladilen A. Trenogin

Parameter continuation method and solvability of nonlinear equations (p.160)

Richard M. Dudley

Differentiability of six nonlinear operators on nonsmooth functions (p.122)

Sect.: 08

Convexity, Banach Spaces, and Isometries, Chair: Themistocles M. Rassias

MA 309 Alexander Ya. Kruger On extremality of set systems (p.137)

Ryszard Grza̧ślewicz

Quasi-interior points for convex sets (p.128)

Themistocles M. Rassias

On the Mazur-Ulam Theorem and the Aleksandrov Problem of Conservative Distances (p.151)

Sect.: 08

Sequence and Function Spaces, Chair: Jan H. Fourie

MA 043 Abdul J. Badiozzaman On a Tauberian Theorem of G.M. Petersen for Abel Summability (p.116)

Detlef Seydel

Matrix Maps of Quasiconvex Null Sequences (p.155)

Jan H. Fourie

On convergence of sections of weakly psummable sequences (p.123)

Sect.: 10

Parabolic Equations, Chair: Nicolay Draganov Kutev

H 111 Helmut Koch Large time regularity of solutions to the porous medium equation (p.210)

Bernd Kawohl

Image processing and anisotropic diffusion (p.208)

Nicolay Draganov Kutev

Maximum and comparison principle for anisotropic diffusion (p.212)

Sect.: 12

Brownian Motion II, Chair: Frank B. Knight

H 107 Viet Ha Hoang Singularly perturbed Dirichlet problems in randomly perforated domains (p.253)

Vesna Perisic

On some energy inequalities (p.265)

Sect.: 12

Statistics for Random Fields, Chair: Alexander Frolovich Taraskin

H 1029 Nikolai Leonenko Parameter identification for stochastic Burgers flows via parabolic rescaling

(p.258)

Nadiya Zinchenko

Asymptotic Properties of Stable Processes and Fields (p.273)

Molisen M. Mohammadzadeh

Spectral Approximation to the Likelihood for an Intrinsic Gaussian Random Field

Sect.: 15

Partial Differential Equations VII, Chair: Ursula van Rienen

H 2033 Maxim A. Olshanskii Iterative method for generalized Stokes problem (p.308)

Chi-Wang Shu

High Order Weighted ENO Schemes for Unstructured Meshes (p.314)

Ursula van Rienen

Efficient Algorithms in Electromagnetic Field Computation (p.316)

Sect.: 17

Optimality Conditions II, Chair: Henrikas Pragarauskas

H 2013 Oleg I. Nikonov Methods of Guaranteed Control Theory in Finance (p.345)

Oleg Vladimirovich Vasiliev

High quality computational methods for optimal control problems (p.353)

Henrikas Pragarauskas

Integrodifferential Bellman equations (p.348)

Poster Session: Tuesday, 25 August, 14:00–15:45

Sect.: 10

Partial Differential Equations

MA 144 Faina S. Berezovskaya About solutions Ltraveling waves of Lienard equation cross-diffusion updatings

(p.198)

Roland Duduchava

Asymptotics of solutions to crack–type and mixed problems (p.202)

Atsuhiko Eida

Harmonic Analysis of Functions (p.203)

Kenji Hasegawa Boundedness of Pseudodifferential Operators (p.206)

Toshio Horiuchi

Removable singularities for quasilinear degenerate elliptic equations with absorption terms (p.206)

Jaan Janno

Inverse problems for determining weakly singular kernels in partial integrodifferential equations (p.207)

Akisato Kubo Global existence of a classical solution to □u − μu + au^{m} = 0 (p.212)

Tokumori Nanbu

On solutions of some nonlinear degenerate parabolic equations (p.215)

Dmitry Shepelsky

Inverse problem for a stratified uniaxial bianisotropic medium (p.220)

Yuming Shi A Foias-Saut type of expansion for a dissipative wave equation (p.221)

Georgi V. Smirnov

A Periodic Knitting Problem for the Helmholtz Equation (p.222)

Anna Surkova

On the localization of the surface states for some class of the discrete Schrödinger oparators with surface potential

Kiyoshi Takeuchi Higher-codimensional boundary value problems for D-modules (p.222)

Jan Turo

Solution of first order stochastic functional partial differential equations via integral contractors (p.223)

Leonid Volevich

Stable and unstable manifolds for elliptic equations with parameter (p.223)

Poster Session: Tuesday, 25 August, 16:15–18:00

Sect.: 09

Ordinary Differential Equations and Dynamical Systems

MA 144 Alina Gleska Comparison theorems for the asymptotic behavior of solutions of difference equations (p.176)

Renato Iturriaga

Lagrangian Graphs, Minimizing Measures, and Mañe’s Critical Values (p.178)

Géza Makay

Marachkov type stability results for functional differential equations (p.181)

Malgorzata Migda On the asymptotic behavior of solutions of nonlinear difference equation (p.183)

Norihiro Misaki

Application of Apparent Monodromy Groups (p.183)

Hiroe Oka

Connectivity of iso-topological entropy curve for symmetric PL bimodal maps

(p.185)

S. Mary Rees Using Teichmüller Geometry (p.187)

Johannes Riedl

On Crossed Renormalization of Quadratic Polynomials (p.187)

Stefan Siegmund

Normal Forms for Nonautonomous Differential Equations of Carathéodory Type

(p.191)

Jarosław Werbowski Oscillatory solutions of functional equations (p.194)

Inna Sergeevna Yemelyanova

Emmy Noether theorem assimilation in classical dynamics (p.195)

Piotr Zgliczynski

Sharkovskii theorem for multidimensional perturbations of 1-dim maps (p.196)

Special Activities

\underline{19:30 – 21:00 H 104}

International Comparison of Mathematical Studies, University Degrees, and Professional Perspectives

Round table, organized by Ina Kersten (Bielefeld, Germany), among the panelists are: D. Cioranescu, K.-H. Hoffmann, J. Netzel, A. Pelczar, M. Primicerio

Other Events

\underline{14:00 – 15:45 MA 004}

Mathematics in the Third Reich and Racial and Political Persecution A forum with lectures by Joel Lebowitz (Rutgers University, USA): “Victims, Oppressors, Activists, and Bystanders: Scientists’ Response to Racial and Political Persecution” and by Herbert Mehrtens (Universität Braunschweig, Germany): “Mathematics and Mathematicians in Nazi Germany. History and Memory” and a discussion chaired by Jochen Brüning (HU Berlin, Germany).

\underline{16:00 Urania}

Mathematics and Art The best entries by high-school students will be honored, and some math videos will be shown.

\underline{17:30 Urania}

Gero von Randow, Hamburg, Germany Geisterfahrer des Journalismus — Wie kommt die Mathematik in die Zeitung? (in German)

\underline{18:00 Urania}

Movie Seeing is Believing / Hotel Hilbert

\underline{19:30 Urania}

Walter Schachermayer, Wien, Austria The Role of Mathematics in Financial Markets. The lecture is followed by a round-table discussion involving R. Dantschke, H. Föllmer (chair), A. Gruber, G. Stahl.

\underline{20:00 Urania}

VideoMath Festival The “VideoMath Reel”, the 24 best contributions to the VideoMath competition, and the Award Ceremony